Predict the annual income for a female aged 45 with 10 years of education. How much would the predicted income have cha...

  1. Predict the annual income for a female aged 45 with 10 years of education. How much would the predicted income have changed for a male? [3.5 points]

Plot the standardized residuals against predicted income,  from regression in part (a). Check for outliers and explain whether the residual plot supports the assumptions about Ɛ. What is your conclusion? Submit the graph to earn full points.

EDUCATION AGE GENDER INCOME (in $1000)
12 60 female 6.5
16 39 male 120
16 33 female 21.75
12 51 male 82.5
16 42 female 55
14 20 male 7.5
14 57 male 37.5
13 61 female 5.5
16 31 male 9
12 30 male 37.5
14 68 female 13.75
16 50 male 32.5
12 27 male 0.5
16 30 male 55
18 65 female 55
19 36 male 67.5
12 22 male 21.75
6 35 male 21.75
12 67 female 9
12 48 male 23.75
12 48 female 45
15 42 male 120
14 61 female 37.5
13 34 male 82.5
17 53 male 82.5
12 39 male 67.5
16 61 male 175
18 34 male 100
12 39 female 45
14 32 male 37.5
16 54 female 45
14 55 female 13.75
14 62 male 32.5
6 39 male 16.25
12 30 female 32.5
12 35 female 16.25
16 55 male 175
17 43 male 175
16 71 male 100
16 55 male 100
14 68 female 45
11 47 male 82.5
16 30 male 55
16 38 female 100
16 41 female 45
20 62 female 120
20 49 male 67.5
16 52 female 100
16 52 male 82.5
14 33 male 82.5
Answer #1

Answer:

  1. Predict the annual income for a female aged 45 with 10 years of education. How much would the predicted income have changed for a male? [3.5 points]

MINITAB used.

The variable gender is coded as female = 1 and male =0.

Regression Equation

INCOME (in $1000)

=

-61.9 + 0.691 AGE - 34.2 Gender + 7.13 EDUCATION

When age =0, female and 10 years education,

predicted income (in 1000) = -61.9+0.691*45-34.2*1+7.13*10 =6.295

or $6295

When age =0, male and 10 years education,

predicted income (in 1000) = -61.9+0.691*45-34.2*0+7.13*10 =40.495

or $40495.

For male predicted value increases by $34200.

Regression Analysis: INCOME (in $1000) versus AGE, ... r, EDUCATION

Analysis of Variance

Source

DF

Adj SS

Adj MS

F-Value

P-Value

Regression

3

35265

11755

8.58

0.000

AGE

1

3544

3544

2.59

0.115

Gender

1

12136

12136

8.85

0.005

EDUCATION

1

19171

19171

13.99

0.001

Error

46

63046

1371

Lack-of-Fit

43

59746

1389

1.26

0.495

Pure Error

3

3301

1100

Total

49

98312

Model Summary

S

R-sq

R-sq(adj)

R-sq(pred)

37.0213

35.87%

31.69%

25.70%

Coefficients

Term

Coef

SE Coef

T-Value

P-Value

VIF

Constant

-61.9

30.5

-2.03

0.048

AGE

0.691

0.430

1.61

0.115

1.18

Gender

-34.2

11.5

-2.98

0.005

1.14

EDUCATION

7.13

1.91

3.74

0.001

1.04

Regression Equation

INCOME (in $1000)

=

-61.9 + 0.691 AGE - 34.2 Gender + 7.13 EDUCATION

Fits and Diagnostics for Unusual Observations

Obs

INCOME
(in
$1000)

Fit

Resid

Std
Resid

27

175.0

94.4

80.6

2.27

R

37

175.0

90.2

84.8

2.36

R

38

175.0

89.0

86.0

2.38

R

R Large residual

Plot the standardized residuals against predicted income,  from regression in part (a). Check for outliers and explain whether the residual plot supports the assumptions about Ɛ. What is your conclusion? Submit the graph to earn full points.

The residuals fan out from left to right rather than exhibiting a consistent spread around the residual = 0 line. The standardized residual vs. fits plot suggests that the error variances are not equal.

Scatterplot of SRES vs FITS SRES o 20 40 60 FITS 80 100 120